# Chypnosis: Stealthy Secret Extraction using Undervolting-based Static Side-channel Attacks

### v2

@@ 3. Threat Model
In our threat model, we assume that the adversary has physical access to the target device. We consider profiling (template) attack scenarios in which the adversary can profile a training device without countermeasures before launching an attack on the actual target. During the attack phase, we assume that all detection- and response-based countermeasures described in Sect. 2.3 are active. The adversary does not have access to the system’s clock source; however, to read the contents of registers at a specific clock cycle, she must halt the clock. We also assume the adversary can access and manipulate the IC’s core voltage supply rails. Moreover, the adversary can capture snapshots of the hardware state using techniques such as LLSI or IA and subsequently recover the values stored in registers. Additionally, we assume the adversary has some knowledge of the system’s clock frequency to estimate the brownout voltage thresholds.
@@
To understand how an adversary may benefit from such an attack in the real world, we can consider the following examples. One example would be the decryption core on FPGAs or microcontrollers/microprocessors, which is programmed with a cryptographic key. Such decryption engines can be used, for example, to decrypt incoming communication traffic or bitstream/firmware of the device. By extracting the secret key, the adversary can access the plaintext communication or clone, reverse engineer, and tamper with the design contained in the bitstream/firmware. Moreover, if the same key is used for multiple ICs in the field, the attacker can break the security of other ICs having the same key.
4. EXPERIMENTAL SETUP
4.1. Devices under test
For IA, we used a NewAE CW305 board (NAE-CW305) (NewAE, 2025b) as our Device Under Test (DUT), which has an AMD/Xilinx Artix-7 FPGA (part number XC7A100T-FTG256) manufactured with 28 nm technology. For LLSI, we used a modified ChipWhisperer CW310 Bergen Board (NewAE, 2025a) as our DUT, which has an AMD/Xilinx Kintex-7 flip-chip FPGA (part number XC7K410T-FBG676), also manufactured with 28 nm technology. We selected this board because FPGA is packaged in a flip-chip package, allowing access to the backside silicon for LLSI. In addition, both the CW305 and CW310 provide direct access to the FPGA’s PDN, which was the primary reason for selecting this board. Notably, we focused on the $V_{CCINT}$ power rail, as it directly powers the FPGA’s core logic and registers. Fig. 4 depicts the high-level diagram of our experimental setup for both attacks, where the measurements are carried out with an external controller.
4.2. Optical Setup
To perform optical probing and photon emission analysis, we used a Hamamatsu PHEMOS-X FA microscope. The system is equipped with a 1.3 µm High-power Incoherent Light source (HIL) and a cooled InGaAs camera operating at –70C. It supports objective lenses with 5x/0.14 NA, 20x/0.4 NA, and 50/0.76 NA magnifications, along with additional scanner zoom levels of 2x, 4x, and 8x.
In LVI/LLSI mode, the laser is scanned across the surface of the device using galvanometric mirrors. The reflected light is separated via semi-transparent mirrors and directed to a photodetector. Its output is passed through a bandpass filter tuned to the frequency of interest. The measured signal amplitude at each scan location is mapped to its corresponding spatial position, forming a grayscale-encoded 2D image.
Photon emission is another failure analysis technique that captures weak light emissions from integrated circuits using a highly sensitive InGaAs camera with long exposure times (typically 5 s or more). When current flows through a P-N junction, it can emit a small number of photons due to energy level transitions, which the cooled InGaAs sensor can detect. For this experiment, we used the PHEMOS-X InGaAs camera with a 20x lens to observe photon emission from a 15-stage ring oscillator implemented on the CW310. This technique was employed to monitor dynamic internal signals, which we expect to become static upon entering hibernation.
4.3. IA Attack Setup
Figure 4. LLSI and IA Setup. Blue components are used only in IA, red only in LLSI, and the rest are common to both setups.
To control the state of the target FPGA on the CW305 board during IA, we used a NewAE CW-Lite board (NewAE, 2025). It facilitates serial communication with the DUT and acts as an intermediate controller for transferring plaintext and receiving ciphertext from the target IC during profiling. The $V_{CCINT}$ voltage on the CW305 can be controlled through USB and an onboard programmable voltage regulator using Python APIs. We used a Keysight ENA Network Analyzer E5080A (Keysight, 2025) for our measurements, which supports RF/microwave scattering analysis across frequencies ranging from 9 kHz to 6 GHz. To ensure reliable signal transmission, we used Minicircuit CBL-2FT-SMNM+ shielded characterization cables (MiniCircuits, 2025), which are also rated for operation within the same frequency range. The hardware designs were implemented in Verilog and synthesized using AMD/Xilinx Vivado Design Suite version 2023.2. The analyzer system, which manages both data preparation and analysis, is a computer running Ubuntu 20.04 LTS. We developed Python scripts to automate the experimental process, including input scheduling, measurement coordination, and attack execution. Instrument control and statistical analysis were performed using the PyVISA and SciPy.stats libraries, respectively.
The measurement process begins loading the the desired bitstream to target IC. Arbitrary input data (e.g., plaintexts and keys) are generated by the Analyzer System and sent to the Controller, see Fig. 4. The controller transmits this data to the target IC via a serial interface. At a designated timestamp, the system drops the voltage to pause the clock; then, the controller triggers the VNA to capture a measurement. The VNA then collects the trace and returns it to the Analyzer System.
The VNA measurement parameters in our analysis are determined experimentally. The intermediate frequency (IF) bandwidth is set to 500 Hz, and the averaging factor is configured as $N_{\text{Avg}}=400$ to reduce the measurement noise floor. We configure the VNA to perform a single-port measurement of $S_{11}$, the reflection scattering parameter, which provides a linear estimation of the impedance profile. Although both amplitude and phase values of the scattering parameters are recorded, we use only the phase component in our impedance analysis due to its superior noise resilience (Mosavirik et al., 2023a). All measurements are performed differentially using a reference program, and the VNA output power is set to 10 dBm. Upon completion of the computation on the target IC, the controller receives the output(i.e., ciphertexts) and forwards it back to the Analyzer System over a serial connection for verification. Fig. 5(a) shows our IA setup.
4.4. LLSI Attack Setup
Figure 5. (a) ChipWhisperer CW305 under impedance attack by the Keysight VNA. (b) ChipWhisperer CW310 Bergen Board inside HAMAMATSU PHEMOS-X microscope chamber modified for LLSI with a custom Arduino board plugged into the PMOD connectors.
For LLSI, we used the same high-level procedure as shown in Fig. 4. We desoldered the bridge between TP20 and TP19 on the CW310, which cuts off $V_{CCINT}$ from the onboard voltage regulator. In its place is one channel of a BK Precision 9130 power supply connected to the SMA connector at J3 (VCCINT_SHUNTLO). This is because the onboard voltage regulator cannot supply a DC voltage that is low enough to hibernate the FPGA.
@@ 4.4. LLSI Attack Setup
We programmed the DUT with registers clocked at 10 MHz with an on-chip MMCM. We can independently set each register to a constant 1, constant 0, or flip with every clock cycle via a USB serial port through an Arduino(i.e., the controller) connected to the PMOD connectors. Although we designed a custom board specifically to fit in the PMOD connectors for convenience (Mitard, 2024), any Arduino board with 3.3V logic would work for this purpose. Fig. 5(b) shows the CW310 Bergen Board and our custom Arduino inside the HAMAMATSU PHEMOS-X microscope chamber.
5. Circuit Operation During Hibernation
5.1. Hibernation Voltage Characterization
5.1.1. Undervolting Voltage-Frequency Scan
@@
To characterize the resilience of the FPGA target under voltage stress, we developed a systematic Hibernation Scan methodology that explores the functional limits of an FPGA’s internal logic as supply voltage decreases across a range of operating frequencies. This method reveals the threshold conditions under which the FPGA ceases to reliably perform core operations such as register assignment and clock-driven state progression. The procedure is executed through a software controller that communicates with an on-chip UART-based hardware test module implemented in RTL within the target FPGA DUT.

### v5

@@ 3 Threat Model
In our threat model, we assume that the adversary has physical access to the target device. We consider profiling (template) attack scenarios in which the adversary can profile a training device before launching an attack on the actual target. During the attack phase, we assume that all detection- and response-based countermeasures described in Section 2.1 are active. We further assume that the security sensors’ internals may be unknown to the attacker. The adversary does not have access to the system’s clock source; however, to read the contents of registers at a specific clock cycle, she must halt the clock. We also assume the adversary can access and manipulate the IC’s core voltage supply rails and remove the decoupling capacitors on the printed circuit board (PCB). Such tampering attempts require some knowledge of the PCB schematic, which can be obtained through documentation, visual analysis, or multimeter testing. Moreover, the adversary can capture snapshots of the hardware state using techniques such as LLSI or IA and subsequently recover the values stored in registers. For IA, the attacker must have access to VNAs and function generators, which are standard RF instruments. For LLSI, failure-analysis equipment can typically be rented for a few hundred dollars per hour rather than purchased outright. For secret extraction, we consider a template attack threat model: (i) LLSI requires localization of target registers, achievable through reverse-engineering or prior design knowledge; (ii) IA requires knowledge of the cryptographic algorithm, masking scheme, and key size. We also assume the adversary has some knowledge of the system’s clock frequency to estimate the brownout voltage thresholds.
@@
To understand how an adversary might benefit from such an attack in the real world, we can consider the following examples. One example would be the decryption core on FPGAs or microcontrollers/microprocessors, which is programmed with a cryptographic key. Such decryption engines can be used, for example, to decrypt the device’s bitstream or firmware. By extracting the secret key, the adversary can clone, reverse-engineer, and tamper with the design contained in the bitstream or firmware. Moreover, if the same key is used for multiple ICs in the field, the attacker can compromise the security of other ICs that use the same key.
4 Experimental Setup
4.1 Devices under Test (DUT)
To test undervolting effects on the clock and voltage sensors, we used two chip families: AMD/Xilinx 7 Series FPGAs and Microchip PolarFire SoC FPGAs, both manufactured using 28 nm technology. While the latter has dedicated hard IP (ASIC) clock and voltage sensors [27], the former has only a hard IP voltage sensor [48], and the clock sensor should be configured as a soft IP (e.g.,  [13]).
4.1.1 AMD/Xilinx FPGAs
We used NewAE CW305 boards [38], which have an AMD Artix-7 FPGA (part number XC7A100T-FTG256). We also used a ChipWhisperer CW310 Bergen Board [39], which has an AMD Kintex-7 flip-chip FPGA (part number XC7K410T-FBG676). We selected these boards primarily because they both provide direct access to the FPGA’s PDN. Furthermore, the OpenTitan design is compatible with the CW310 Bergen Board. We focused on the $V_{\mathit{CCINT}}$ power rail, as it directly powers the FPGA’s core logic and registers. Moreover, these boards do not contain any decoupling capacitors on the core voltage PDN, making a rapid voltage drop feasible.
4.1.2 Microchip FPGA SoCs
@@
4.2 Optical Setup
To perform optical probing and photon emission analysis, we used a Hamamatsu PHEMOS-X FA microscope [19]. The system is equipped with a 1.3 µm High-power Incoherent Light source (HIL) for optical probing. It supports objective lenses with 5x/0.14 NA, 20x/0.4 NA, and 50/0.76 NA magnifications, along with additional scanner zoom up to 8x.
In LVI/LLSI mode, the laser is scanned across the device’s surface using galvanometric mirrors. The reflected light is separated via semi-transparent mirrors and directed to a photodetector. Its output is passed through a bandpass filter tuned to the frequency of interest. The measured signal amplitude at each scan location is mapped to its corresponding spatial position, forming a grayscale-encoded 2D image.
Photon emission is another backside failure analysis technique that captures weak light emissions from the chip’s transistors with long exposure times (typically 5 s or more). When current flows through a P-N junction, it can emit a small number of photons due to energy level transitions. For this experiment, we used the PHEMOS-X InGaAs camera operating at –70C with a 20x lens to observe photon emission from a 15-stage ring oscillator implemented on the CW310. This technique was employed to monitor dynamic internal signals, which we expect to become static upon entering hibernation.
4.3 Side-channel Attack Setups
We used the CW305 and CW310 boards for all side-channel attacks. Figure 4 depicts the high-level diagram of our experimental setup for IA and LLSI, where the measurements are carried out with an external controller.
4.3.1 LLSI Attack Setup
For LLSI, we used the same high-level procedure as shown in Figure 4. We desoldered the bridge between TP20 and TP19 on the CW310, which cuts off $V_{\mathit{CCINT}}$ from the onboard voltage regulator. In its place, we connect one channel of a BK Precision 9130 power supply to the SMA connector at J3 (VCCINT_SHUNTLO). This is because the onboard voltage regulator cannot supply a DC voltage that is low enough to hibernate the FPGA.
[…]
@@ 4.3.1 LLSI Attack Setup
We programmed the DUT with registers clocked at 10 MHz with an on-chip MMCM. We can independently set each register to a constant 1, a constant 0, or flip with every clock cycle via a USB serial port through an Arduino (i.e., the controller) connected to the PMOD connectors.
4.3.2 IA Attack Setup
Figure 4: LLSI and IA Setup. Blue parts are used only in IA, red only in LLSI, and the rest are common to both.
To control the state of the target FPGA on the CW305 board during IA, we used a NewAE CW-Lite board [40]. It facilitates serial communication with the DUT and acts as an intermediate controller for transferring plaintext and receiving ciphertext from the target IC during profiling. The $V_{\mathit{CCINT}}$ voltage on the CW305 can be controlled through USB and an onboard programmable voltage regulator using Python APIs. We used a Keysight ENA Network Analyzer E5080A [21] for our measurements, which supports RF/microwave scattering analysis across frequencies ranging from 9 kHz to 6 GHz. To ensure reliable signal transmission, we used Minicircuit CBL-2FT-SMNM+ shielded characterization cables [30].
After loading a bitstream into the target FPGA, arbitrary input data (e.g., plaintexts and keys) are generated by the Analyzer System and sent to the Controller, as shown in Figure 4. The controller transmits this data to the target IC via a serial interface. At a designated timestamp, the system drops the voltage to pause the clock; then, the controller triggers the VNA to capture a measurement. The VNA then collects the trace and returns it to the Analyzer System.
The VNA measurement parameters in our analysis are determined experimentally. The intermediate frequency (IF) bandwidth is set to 500 Hz, and the averaging factor is configured as $N_{\text{Avg}}=400$ to reduce the measurement noise. We configure the VNA to perform a single-port measurement of $S_{11}$, the reflection scattering parameter, which provides a linear estimation of the impedance profile. Although both amplitude and phase values of the scattering parameters are recorded, we use only the phase component in our impedance analysis due to its noise resiliency [36]. All measurements are performed differentially using a reference program, and the VNA output power is set to 10 dBm. Upon completion of the computation on the target IC, the controller receives the ciphertexts and forwards them back to the Analyzer System over a serial connection for verification.
5 Circuit Operation During Hibernation
5.1 Hibernation Voltage Characterization
5.1.1 Undervolting Voltage-Frequency Scan
@@
To characterize the resilience of FPGA targets under voltage stress, we developed a systematic Hibernation Scan methodology that explores the functional limits of an FPGA’s internal logic as supply voltage decreases across a range of clock frequencies. This method reveals the threshold conditions under which the FPGA ceases to reliably perform core operations such as register assignment and clock-driven state transitions. The procedure is executed through a software controller that communicates with an on-chip UART-based hardware test module implemented in RTL within the target FPGA DUT. Crucially, this UART module is on a separate voltage rail to the FPGA internal logic, so it remains functional during our undervolting experiments.